OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

answer is d again like pre problem slope should be +1 y-(-2)=1(x+4) y=x+2

OpenStudy (zehanz):

Product of slopes of two perpendicular lines is -1, so, like shining said, slope of new line = 1. The general formula of a line with slope m through point (a, b) is: y-b=m(x-a). So all you have to do now is fill in m=1, a=-4, b=-2, this leads to: y--2=1(x--4) <=> y+2=x+4 <=> y =x+2, so it's answer d.